OpenStudy (anonymous):

Tammy, working alone, could clean one-half the area in one hour. Her partner Melissa, working alone could clean one-third of the area in one hour. How long would it take them to clean the area if they worked together?

jimthompson5910 (jim_thompson5910):

tammy's rate is 1/2 a job per hour Melissa's rate is 1/3 of a job per hour add the two fractions to get ???

OpenStudy (anonymous):

you get 5/6

OpenStudy (anonymous):

but then i want to know how long it would take them to clean up the whole area

jimthompson5910 (jim_thompson5910):

so their combined rate is 5/6 of a job per hour

jimthompson5910 (jim_thompson5910):

(rate)*(time) = 1 full job (5/6 job per hr)*(x hours) = 1 full job (5/6)*x = 1 x = ??

OpenStudy (anonymous):

do i divide the job per hr from 1 full job..?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

Multiply both sides by the reciprocal of 5/6